how do you find the value of variables?
example:
if angle a=x degrees, angle b=(x+40) degrees, angle c=(x-50) degrees
angle A is supplementary to angle B and complimentary to angle C

Answer by stanbon(75887)

if angle a=x degrees, angle b=(x+40) degrees, angle c=(x-50) degrees
angle A is supplementary to angle B and complimentary to angle C
-------------------------------
Equations:
x + x+40 = 180
------
2x + 40 = 180
2x = 140
x = 70
==============
x + x-50 = 90
----
2x = 140
x = 70
=========
A = x = 70 degrees
B = x+40 = 110 degrees
C = x-50 = 20 degrees
=========================
Cheers,
Stan H.
==============